# Expressive Limits of Quantum Reservoir Computing

### v1

@@ III Limits of expressivity given by the input encoding
The quantum circuit with a single-qubit encoding via a rotation gate has access to a single frequency as presented in Ref. [24], which leads to a simple sine function as eigentask. Furthermore, the circuit can express a function that is independent of the input, i.e. a constant. This results in a REC of two, as fulfilled by the product state. The same circuit with one rotational input gate and an additional entangling layer has still access to a single frequency. The entangling layer enables the ‘distribution’ of the input information to other degrees of freedom. In this way, the circuit can construct an additional orthogonal function of the same frequency, which in this case is the cosine that appears in Fig. 4. For every rotational input gate that is added, the accessible frequency spectrum is extended by one frequency, increasing the number of eigentasks by two if an an entangling layer is being used. Hence, the input-gate-number dependent upper bound (Eq. (15)) is perfectly replicated by our REC calculations shown in Fig. 5.
@@
An exception to this is the case of using eight encoding gates for the input, where Eq. (15) predicts a value of $C_{\text{T}}=17$. The reason for this is that this the REC value is not only limited by the number of input gates, but also by the number of the measured degrees of freedom as elaborated in Ref. [25]. The number of measured degrees of freedom is 16 in our setup due to the POVM on four qubits, restricting the REC to the same value.
@@
Fig. 5 also shows the expressivity in dependence of the number of input gates in product state systems. These systems feature a reduced expressivity, because the information is injected at the level of the individual qubits, in which case the output can only be composed as a superposition of an incomplete set of eigenvalues of the total many-particle system.
@@
Reaching the upper bound in Eq. 15 is only possible if the entangling layer allows for the distribution of the input information into different degrees of freedom of the system. Consequently a layer of CNOT gates possesses the same REC value as the product state [25]. An important aspect that we have not yet considered is the presence of noise, which is always present in real-world implementations. Finite noise-to-signal ratios (NSRs) prevent reaching this upper bound, leading to the question, which ansatz circuits maximize the expressivity under such conditions. In the following two sections, we investigate the expressivity of systems where we use the time-evolution of the TFIM as reservoir in the presence of shot noise and compare it with different ansatz circuits.
IV Resolvable expressive capacity in QRC
Up to now, we used as reservoir $U_{\text{res}}$ only gate-based ansätze. Next, we apply the expressivity analysis to the field of physical QRC. Meaning that by using the methods from Section II, we exchange $U_{\text{res}}$ by the unitary time-dynamics of a physical quantum system, chosen to be the TFIM.
The setup we consider is the one depicted in Fig. 1, but for the input encoding, we now specify two different scenarios that are illustrated in Fig. 6.
Figure 6: QRC scheme for determining the REC, where two cases are considered: First case is without the random rotations (symmetric system) and the second case includes the random rotations to break the symmetry.
In the first case, the input encoding layer is directly followed by $U_{\text{res}}$, i.e. the dynamics of the TFIM as given in Eq. 8. The second case adds a layer of random single qubit rotations after each input layer to break the symmetry. This layer has been shown to be beneficial, as the presence of symmetries in the system limits the attainable expressivity below its theoretical maximum [25]. Since in the TFIM all qubits are treated equally, our system possesses an intrinsic symmetry due to the Hamiltonian being not only Hermitian but also symmetric. In an experimental setup of the physical system, the breaking of symmetry might already occur due to the presence of noise of such an open quantum systen.
Figure 7: REC values of the TFIM in dependence of the input encoding.
The statistical nature of the quantum-measurement processes is not accounted for at this stage by assuming exact expectation values. We will explicitly consider the impact of the measurement statistics in Section V.
The REC that persists under these circumstances is shown in Fig. 7 in dependence on the number of input encodings. For the first three encoding gates, the QRC based on the TFIM behaves similarly to the quantum circuit with an entangling ansatz. No difference between the symmetric and the broken-symmetry system is observed, because the input is injected on only three out of four qubits, thereby destroying the underlying symmetry. However, adding a fourth encoding gate introduces a discrepancy between both systems, because the input is injected symmetrically on all qubits. Increasing the number of input gates further, it becomes apparent that the broken-symmetry system exhibits the maximal achievable REC values following our upper bound in Eq. 15, meaning that the time evolution of the TFIM is maximally expressive in a theoretical ideal and symmetry broken scenario. In a system with symmetries, the expressivity is reduced as already mentioned in Ref. [25]. In the extreme case of eight encoding gates, we find an attainable REC of $C_{\text{T}}=8$, caused on the one hand, by the symmetric Hamiltonian, and, on the other hand, also by the symmetric input encoding. One should point out that in an experimental realization, depending on the underlying hardware platform, perfectly symmetric systems will not likely be realized due to inhomogeneity and dissipation processes.
V Gate-based ansätze for a QRC scheme
@@
In this section we aim for a more realistic calculation of the expressivity and consider the impact of shot noise when performing a finite number of measurements. This will generally reduce the REC due to the non-zero noise-to-signal ratios (NSRs) of the eigentasks. The NSRs depend on the underlying system, in particular on the used reservoir.
@@
Although, the idea of RC is to use physical systems as reservoir implementation rather than programmable (quantum) computers, the gate-based approach offers an opportunity to investigate the general properties of quantum systems in RC. Furthermore, the gate-based architecture could offer a more approachable platform for RC in the NISQ era. Here, we analyse the use of the different ansatz circuits from Sim et al. [35] (refer to Fig. 10) as reservoirs.
@@
Figure 8: Blue: REC for different ansatz circuits. Values are calculated by randomly sampling circuits parameters and averaging over 200 realizations. Green: REC values of the circuits with optimized, rather than randomly sampled, parameters. Parameters of the ansatz circuits have been optimized for a maximal expressivity. Dashed gray line indicates the expressivity of the non-optimized TFIM optimized over 50 reservoir realizations. Shaded gray area indicates the standard deviation.
@@
The REC values for all considered circuits are shown in Fig. 8. As expected, a product state, as realized by circuit 1, has the lowest expressivity. Circuit 2, which produces an entangled state via a single layer of CNOT gates, possesses the same expressivity. Again, this shows that entanglement alone is not the reason for a circuit to be expressive. Circuits 3, 16 and 18 also have a low expressivity, which could be explained by the fact that, in these circuits, the controlled rotations address the same degrees of freedom as the single qubit rotations prior to them. Apart from that, the circuits 13 and 14 with controlled rotations perform quite well in general. For the expressivity, it seems to be beneficial if ansatz circuits have alternating parts between entangling gates and single-qubit rotations. Especially circuit 15, which uses solely CNOT gates for entanglement generation, shows large improvement compared to circuit 3. Also the high-depth circuits 5 and 6, in which entangling gates and single-qubit rotations are not mixed, perform worse than circuits 13 and 14.
@@
The shown results demonstrate that, in contrast to classical reservoirs and possibly against intuition, QRCs get their expressivity not from the reservoir. Once, the reservoir introduces sufficient complexity for achieving maximal expressivity as allowed due to the input encoding, further increasing the size of the reservoir or its complexity cannot improve the expressivity. The expressivity is primarily determined by two factors: firstly by the way the input is injected as shown here, and secondly by the number of readouts that are preformed [25]. The reservoir can only change the number of functions that are accessible, but the source for expressivity lies in the input injection. To be precise, the set of non-linear orthogonal functions the reservoir has access to is independent of the reservoir, since the time evolution in quantum mechanics is inherently linear. Non-linearity can only be introduced by the encoding gates.
@@
Before concluding, we address the expressivity of the physical QRC that is also shown in Fig. 8. Interestingly, the TFIM has the lowest REC value among all tested reservoirs, performing even worse than the product-state circuit. We attribute this result to the specific dynamics produced by the time evolution of the TFIM, leading to output functions with high NSRs, thus lowering the expressivity. Besides the higher expressivity of the ansatz circuits, the gate-based approach to QRC has the additional advantage over pyhsical reservoirs that their parameters can be precisely tuned and, therefore, be used for optimization. To do so, we employ an approach similar to parameterized quantum circuits in QML by performing an optimization step of the parameters of all ansatz circuits shown in Fig. 10 to maximize the attainable expressivity. The REC values for the optimized circuits are shown in green in Fig. 8. The results indicate that this approach generally improves the performance, although the increase in expressivity varies significantly among the individual quantum circuits. While the REC values of the worst performing circuits get a large increase after optimization, circuits 14 and 15 improve only slightly. Generally, after optimization there are less differences in the performance of the quantum circuits indicating that optimization is worthwhile in any case.
@@
An open question remains how the noise-to-signal ratios of the eigentasks are affected by the specific choice of the circuit, which will be the topic of future investigations.
VI Conclusion
@@
The central conclusion of our work is the counter-intuitive fact that in physical reservoir computing, there is a fundamental difference between classical and quantum hardware platforms: While classical RC draws its capabilities from the complex non-linear dynamics that is inherent to the reservoir, the size and connectivity of a quantum reservoir does not determine its expressivity, which serves as a measure for the complexity of the output that can be generated from the input data.
@@
To reveal this fact, we have translated the QRC approach into the language of circuit-based QML, establishing a fundamental connection on an abstract level between these two very different QML paradigms. This analogy enables one to compare the expressivity of the TFIM as a representative for a physical QRC with a selection of different gate-based low-depth ansatz circuits. We find all of these ansatz circuits to outperform the TFIM with the additional benefit that they can be optimized via their parametrization, which is generally not possible in a physical reservoir. We explicitly show that the upper bound for the number of linear independent functions a reservoir can express is limited by the number of input encoding gates. Considering the inherent linearity of quantum mechanics, it is impossible to increase the available number of linear independent functions by increasing the complexity or the size of the quantum reservoir. In other words, the available function space is solely determined by the number of input encoding gates, while the reservoir topology merely influences the number of orthogonal functions that are accessible.
The insight we provide carries an important message into the field of physical QML and QRC, namely that the exponentially large state space one finds in quantum mechanics cannot be used as an advantage of quantum over classical RC. Harnessing this quantum advantage requires the input encoding to address the available degrees of freedom that, otherweise, remain unaccessible for computing. While this problem could be mitigated at the input-encoding stage, doing so will require precise control over the input encoding, which is generally not the idea behind physical QRC and is further limited by the NISQ hardware that is currently available.
@@ Acknowledgements
The authors would like to thank Frederik Lohof for many useful discussions. This work has been supported by the Quantum Computing Initiative of the DLR via the Quantum Fellowship Programme. We also acknowledge funding from the German Research Foundation (DFG) via the project Gi1121/6-1.
Appendix A Transpilation of the time-evolution of the TFIM
Simulation of the TFIM on a gate-based quantum computer requires a transpilation of the unitary time-evolution of the Ising system into the native gate set of the used quantum computer. As example, we transpiled the time-evolution operator, as given by Eq. (5), into the native gate-set of the IBM kyiv device. We performed the transpilation for a single multiplexing step of $\Delta_{t}/V$. The transpiled circuit is shown in Fig. 9.
@@ Appendix B Ansatz circuits
For convenience, we show the ansatz circuits used by Sim et al. [35] in Fig. 10.

### v2

@@ III Limits of expressivity given by the input encoding
The expressivity results of the product system and the one with this entangling ansatz are shown in Fig. 5. The quantum circuit with a single-qubit encoding via a rotation gate, as defined by Eq. (13), has access to a single frequency as presented in Ref. [28], which leads to a simple sine function as eigentask. Furthermore, the circuit can express a function that is independent of the input, i.e. a constant. This results in a REC of two, as fulfilled by the product state. The same circuit with one rotational input gate and an additional entangling layer has still access to a single frequency. The entangling layer enables the ‘distribution’ of the input information to other degrees of freedom. In this way, the circuit can construct an additional orthogonal function of the same frequency, which in this case is the cosine that appears in Fig. 6. For every rotational input gate that is added, the accessible frequency spectrum is extended by one frequency, increasing the number of eigentasks by two if an an entangling layer is being used. Hence, the input-gate-number dependent upper bound (Eq. (18)) is perfectly replicated by our REC calculations shown in Fig. 5.
@@
An exception to this is the case of using eight encoding gates for the input injection, where Eq. (18) predicts a value of $C_{\text{T}}=17$. The reason for this is that the REC value is not only limited by the number of input gates, but also by the number of the measured degrees of freedom as elaborated in Ref. [25]. The number of measured degrees of freedom is 16 in our setup due to the POVM on four qubits, restricting the REC to the same value.
@@
Figure 5 also shows the expressivity in dependence of the number of input gates in product state systems. These systems feature a reduced expressivity, because the information is injected at the level of the individual qubits, in which case the output can only be composed as a superposition of an incomplete set of eigenvalues of the total many-particle system.
Reaching the upper bound in Eq. (18) is only possible if the entangling layer allows for the distribution of the input information into different degrees of freedom of the system. Consequently a layer of CNOT gates possesses the same REC value as the product state [25].
Up to now, the reservoir dynamics we investigated were realized by a gate-based ansatz circuit. Now, we apply the expressivity analysis also to the field of physical QRC. By using the methods from Sec. II, we therefore exchange $U_{\text{res}}$ by the unitary time dynamics of a physical quantum system, chosen to be the TFIM. Hence, the setup is the same as depicted in Fig. 2 with $U_{\text{res}}=\text{e}^{-\text{i}H\Delta_{t}}$. The REC analysis reveals that the TFIM has the same dependence on the number of input gates as the gate-based ansatz circuit used before, which is also shown in Fig. 5. In a theoretical ideal scenario, the TFIM achieves the maximal possible expressivity. This agrees with our finding that the limits of expressivity in QRC depend on the input encoding procedure, rather than on the underlying system.
@@
An important aspect that we have not yet considered is the presence of noise, which is always present in real-world implementations. Finite signal-to-noise ratios (SNRs) of the eigentasks prevent reaching this upper bound, leading to the question, which ansatz circuits maximize the expressivity under such conditions. In the following section, we investigate the expressivity of systems where we use the time evolution of the TFIM as reservoir in the presence of shot noise and compare it with different ansatz circuits.
IV Gate-based ansätze for a QRC scheme
@@
In this section we aim for a more realistic calculation of the expressivity and consider the impact of shot noise when performing a finite number of measurements to approximate expectation values. This will generally reduce the REC due to the finite signal-to-noise ratios (SNRs) of the eigentasks. The SNRs depend on the underlying system, in particular on the used reservoir.
@@
Although, the idea of RC is to use physical systems as reservoir implementation rather than programmable (quantum) computers, the gate-based approach offers an opportunity to investigate the general properties of quantum systems in RC. Furthermore, the gate-based architecture could offer a more approachable platform for RC in the NISQ era. Here, we analyze the use of the different ansatz circuits from Sim et al. [43] (refer to Fig. 9) as reservoirs.
@@
Figure 7: Expressivity values for different ansatz circuits and the TFIM. For all systems, input is injected via eight single-qubit rotations around randomly sampled axes. Blue: REC for different ansatz circuits. Values are calculated by randomly sampling circuits parameters (20 samples) and randomly sampling the parameters for the arbitrary rotations (30 samples) and averaging over the realizations. Green: REC values of the circuits with optimized, rather than randomly sampled, parameters. Parameters of the ansatz circuits have been optimized for a maximal expressivity. Dashed gray line indicates the expressivity of the non-optimized TFIM averaged over 20 reservoir realizations. Shaded gray area indicates the standard deviation.
@@
We calculated the REC values for all considered circuits using eight input encoding gates around randomly sampled axes. The results are shown in Fig. 7. As expected, a product state, as realized by circuit 1, has among the lowest expressivity. Circuit 2, which produces an entangled state via a single layer of CNOT gates, possesses the same expressivity. Again, this shows that entanglement alone is not the reason for a circuit to be expressive. Circuits 3, 16 and 18 also have a low expressivity, which could be explained by the fact that in these circuits the controlled rotations address the same degrees of freedom as the single-qubit rotations prior to them. Apart from that, the circuits 13 and 14 with controlled rotations perform quite well in general. For the expressivity, it seems to be beneficial if ansatz circuits have alternating parts between entangling gates and single-qubit rotations. Especially circuit 15, which uses solely CNOT gates for entanglement generation, shows a large improvement compared to circuit 3.
The shown results demonstrate that, in contrast to classical reservoirs and possibly against intuition, the ability in QRCs to increase the output function space relies on the input encoding and not on the reservoir dynamics.
@@
Once the reservoir introduces sufficient complexity for reaching the maximal expressivity allowed according to the input encoding, further increasing the size of the reservoir or its complexity cannot improve the expressivity. The expressivity is primarily determined by two factors: firstly by the way the input is injected as shown here, and secondly by the number of readouts that are performed [25]. The reservoir can only change the number of functions that are accessible, but the source for expressivity lies in the input injection. The set of non-linear orthogonal functions the reservoir has access to therefore is independent of the reservoir, since the time evolution in quantum mechanics is inherently linear.
@@
Before concluding, we address the expressivity of the physical QRC that is also shown in Fig. 7 (dashed line shows average REC over 20 reservoir realizations). Interestingly, the TFIM has the lowest mean REC value among all tested reservoirs, performing even worse than the product-state circuit. We attribute this result to the specific dynamics produced by the time evolution of the TFIM, leading to output functions with smaller SNRs, thus lowering the expressivity. Besides the higher expressivity of the ansatz circuits, the gate-based approach to QRC has the additional advantage over physical reservoirs that their parameters can be precisely tuned and, therefore, be used for optimization. To do so, we employ an approach similar to parameterized quantum circuits in QML by performing an optimization step of the parameters of all ansatz circuits shown in Fig. 9 to maximize the attainable expressivity. The REC values for the optimized circuits are shown in green in Fig. 7. The results indicate that this approach generally improves the performance, although the increase in expressivity varies significantly among the individual quantum circuits. While the REC values of the worst performing circuits get a large increase after optimization, circuits 14 and 15 improve only slightly. Generally, after optimization there are less differences in the performance of the quantum circuits indicating that optimization is worthwhile in any case.
@@
An open question remains how the SNRs of the eigentasks are affected by the specific choice of the circuit, which will be the topic of future investigations.
V Conclusion
@@
The central conclusion of our work is the counter-intuitive fact that in physical reservoir computing, there is a fundamental difference between classical and quantum hardware platforms: While classical RC draws its capabilities from the complex non-linear dynamics that is inherent to the reservoir, quantum systems get their expressivity not from the reservoir. Instead, the power to express non-linear functions from the input data (expressivity) relies on the input encoding.
@@
To reveal this fact, we have translated the QRC approach into the language of circuit-based QML, establishing a fundamental connection on an abstract level between these two very different QML paradigms. This analogy enables one to compare the expressivity of the TFIM as a representative for a physical QRC with a selection of different gate-based low-depth ansatz circuits. We find all of these ansatz circuits to outperform the TFIM if shot noise is taken into account, with the additional benefit that they can be optimized via their parametrization, which is generally not possible in a physical reservoir.
Our results challenge the widespread expectation of exponential computational advantage in QRC based solely on Hilbert space dimension, and instead highlight input encoding as the critical bottleneck to expressivity. We anticipate that overcoming this bottleneck will require new input injection paradigms or the use of quantum-native data, opening up new avenues for both experimental and theoretical exploration in the field.
@@ Acknowledgements
The authors would like to thank Frederik Lohof for many useful discussions. This work has been supported by the Quantum Computing Initiative of the German Aerospace Center (DLR) via the Quantum Fellowship Programme. We are grateful for funding from the German Research Foundation (DFG) and the Agence nationale de la recherche (ANR) via the German-French cooperation project PhotonicQRC (DFG: Gi1121/6-1).
Appendix A Transpilation of the time evolution of the TFIM
Simulation of the TFIM on a gate-based quantum computer requires a transpilation of the unitary time evolution of the Ising system into the native gate set of the used quantum computer. As example, we transpiled the time-evolution operator, as given by Eq. (5), into the native gate-set of the IBM kyiv device. We performed the transpilation for a single multiplexing step of $\Delta_{t}/V$. The transpiled circuit is shown in Fig. 8, illustrating the complexity of implementing the inherently simple time evolution of the TFIM on a gate-based device.
@@ Appendix B Ansatz circuits
For convenience, we show the ansatz circuits used by Sim et al. [43] and here in this work in Fig. 9.